OpenStudy (karlaltr):

y=√5-x What is the slope? can I determine domain and range with this? Where is the x-int/y-int?

OpenStudy (wckeller):

This function is still pretty much normal except that it has been moved up √5 units. It can be re-written as \[y=-x+\sqrt{5}\] The equation for a line is y=mx+b The slope is the number before the variable "x". Domain and range are unaffected by this transformation, so it should be easy to find. Because there has been no left or right movement, the x-int has stayed the same. Meanwhile, the y-int has been moved up √5 units.

OpenStudy (karlaltr):

I should have extended the line ugh Im sorry, its supposed to be a sideways parabola so 6-x are under the square root

OpenStudy (karlaltr):

*5-x

OpenStudy (wckeller):

Is this what you meant? \[y=\sqrt{5-x} \]

OpenStudy (wckeller):

The slope of this function involves derivatives, are you sure that you are supposed to find the slope?

OpenStudy (karlaltr):

Well I am just supposed to graph it but Im not sure how to do that without the slope

OpenStudy (wckeller):

You can still graph this function without its slope, just plug in x values and you get the y values. Which allows you to plot points on a graph.

OpenStudy (karlaltr):

so if I plugged 1 into the equation for x, i would fin the corresponding y value, that makes sense idk why i didnt think of that lol

OpenStudy (wckeller):

This function can be re-written as: \[y=\sqrt{-x+5}\] Using the IHOD OVER acronym we know that because the +5 is on the inside of the radical, we are actually moving left 5. The domain of the function would be shifted left 5 units. The range is not affected by the -1 in front of the x variable. The -1 coefficient is on the inside of the radical, the function will be reflected over the Y axis. Thereby not changing what Y values the function reaches. The domain is changed by this coefficient, it is now reversed. Instead of a [0,∞) domain, it is now (-∞,0]

satellite73 (satellite73):

not really

satellite73 (satellite73):

\[f(x)=\sqrt{5-x}\] to find the domain, solve \[5-x\geq 0\] in one step

OpenStudy (wckeller):

*facepalm* I keep messing these up.

OpenStudy (wckeller):

Im sorry @karlaLTR listen to satellite73

OpenStudy (wckeller):

@satellite73 Thank you for fixing my big mistake! :)

OpenStudy (karlaltr):

Thanks for all the help guys :)